OpenStudy (anonymous):

Please Help! Medal and Fan!!!! :) The table below shows the values of y for each value of x: x= 2,3,4,5 y=1,1,-3,-3 Part A: Does the table represent a relation that is a function? Justify your answer by using the values shown in the table. Part B: The function f(h) shown below represents Kylie's total earning from working in a cafe for h hours of work: f(h) = 15h+5 Calculate and interpret the meaning of f(8) Part C: Write and ordered psi to represent the input and output of the function in Part B when Kylie works for r hours.

OpenStudy (anonymous):

@shinalcantara Could you please help me?

OpenStudy (anonymous):

@gorv Could you please help?

OpenStudy (anonymous):

@texaschic101 could you please help me?

OpenStudy (gorv):

what is property of function??

OpenStudy (gorv):

function should be one one

OpenStudy (anonymous):

Its not a function for part A because it repeats right?

OpenStudy (anonymous):

Am I right?

OpenStudy (gorv):

that is one value of y for one value of x

OpenStudy (anonymous):

I have part b but I'm not sure about A and C

OpenStudy (anonymous):

Okay I'm confused :(

OpenStudy (gorv):

you now vertical line test

OpenStudy (anonymous):

? No I'm not on a vertical line test

OpenStudy (gorv):

one x should tell one value of y

OpenStudy (gorv):

it should not tell more thn one value of y

OpenStudy (anonymous):

So it is not a function because Y has more than 1

OpenStudy (gorv):

x=y^2 x=4 mean y=2 or y=-2

OpenStudy (gorv):

if we put y=2 it means x=4 and if y=-2 that also mean x=4 here x=4 alone represnet 2 values of y

OpenStudy (gorv):

so it should not be like this

OpenStudy (gorv):

in given Q there is nothing like this so it is a function

OpenStudy (anonymous):

wait so it is a function...?

OpenStudy (gorv):

yeah

OpenStudy (gorv):

f(h)=15h+5

OpenStudy (gorv):

f(8) means h is replaced by 8

OpenStudy (gorv):

that is value of function when h=8

OpenStudy (anonymous):

I know part b f(h)=15h+5 f(8) = 15(8)+5 f(8)= 120+5 f(8)= 125

OpenStudy (gorv):

f(h)=15h+5 plug in h=8 here and calculate f98)

OpenStudy (gorv):

yeah well done

OpenStudy (anonymous):

:D thanks now part c Im kinda confused on

OpenStudy (gorv):

for r hours means h= r hours so plug in h=r in place of h everywhere

OpenStudy (anonymous):

so like f(h)=15h+5 it would be like f(r)=15r+5 ?

OpenStudy (gorv):

yeah you are right

OpenStudy (anonymous):

Sooo like F(r)= 15r+5 35, 50, 65, 80 (2,35) (3,50) and so on so the answer would be (r, 15r+5)

OpenStudy (anonymous):

^_^ Okay! I think i got it thanks @gorv for all your help i medaled you!!!